OpenStudy (anonymous):

Use the quadratic equation y = x2 – 6x + 9. Where does the vertex of the parabola lie?

OpenStudy (igreen):

Formula to find the x-coordinate of the vertex: \(\sf \dfrac{-b}{2a}\)

OpenStudy (igreen):

Your equation is in the form of: ax^2 + bx + c

OpenStudy (igreen):

So in this case, 'b' is -6 and 'a' is 1.

OpenStudy (igreen):

\(\sf \dfrac{-b}{2a}\) \(\sf \dfrac{-(-6)}{2(1)}\) Simplify

OpenStudy (igreen):

-1 * -6 = ? 2 * 1 = ?